# HOTS3D: Hyper-Spherical Optimal Transport for Semantic Alignment of Text-to-3D Generation

### v1

@@ 3.2 Derivation of SOT solution on hyper-sphere
Here $\mathtt{CVX}(\nu)$ denotes the set of all convex functions in $L^{1}(\nu)$, $g^{c}(\boldsymbol{x})=\inf_{\boldsymbol{y}}(\log(2-\langle\boldsymbol{x},\boldsymbol{y}\rangle)-g(\boldsymbol{y}))$ is the conjugate c-transform of $g(\cdot)$.
According to [67, Theorem 5.26], any optimal pair potential $(g,g^{c})$ in Eq. (6) induce an OT map pushing forward $\nu_{S}$ onto $\mu_{S}$, and $(f,f^{c})$ induce an OT map pushing forward $\mu_{S}$ onto $\nu_{S}$. Specifically, if $c(\boldsymbol{x},\boldsymbol{y})=\frac{1}{2}||\boldsymbol{x}-\boldsymbol{y}||^{2}$, the OT map $T(\boldsymbol{x})=\boldsymbol{x}-\nabla f(\boldsymbol{x})$. When we shift our attention to hyper-spheres, this solution no longer applies. To provide a functional relationship between the OT map and Kontarovich’s potential on hyper-spheres, we introduce some theorems below.
@@
([67, Villani]). Given $\mu$ and $\nu$ on a compact convex domain $\Omega\subset\mathbb{R}^{d}$, there exists an OT plan for the cost $c(\boldsymbol{x},\boldsymbol{y})=h(\boldsymbol{x}-\boldsymbol{y})$, with $h$ strictly convex. It is unique and of the form $(id,T_{\#})\mu$ (id: identity map), provided that $\mu$ is absolutely continuous and $\partial\Omega$ is negligible. Moreover, there exists a Kantorovich’s potential $\varphi$, and the OT map $T$ can be represented as follows:
@@
This theorem was first explained by Villani [67] and later organized by Lei et al. [68, 69]. Based on it, we further provide an explicit functional relationship between the OT map and Kantorovich’s potential $\varphi$, in the case of hyper-spheres.
@@
Given two probability measures $\mu_{S}$ and $\nu_{S}$ which are defined on $\boldsymbol{X}\subset\mathbb{S}^{d}$ and $\boldsymbol{Y}\subset\mathbb{S}^{d}$, respectively. We have
@@
[…]
According to Theorem 3.2, there exists a potential function $f$ such that $T_{f}:\boldsymbol{X}\to\boldsymbol{Y}$ in Eq. (15) satisfies (4). This indicates that there exists an optimal pair $(f_{0},g_{0})$ achieving the infimum and supremum respectively, and $T_{f_{0}}$ is the OT map from $\mu_{S}$ to $\nu_{S}$. In other words, $T_{f_{0}}$ maps $\boldsymbol{X}$ to $\boldsymbol{Y}$ in a measure preserving manner, satisfying $T(\boldsymbol{X})=\boldsymbol{Y}$. Thus, the solution for Eq. (6) is also a solution for Eq. (14). Based on this, the next pertinent question is how to construct a suitable $f$ and $g$, and how to effectively solve this optimization problem.
Algorithm implementation for SOT learning. Inspired by Makkuva et al. [70], we instantiated the Kantorovich potential $f$ and $g$ with ICNN [26], where ICNNs are a specific parametric class of convex functions with rich enough representation. The network architecture of potential $f$ and $g$ is shown in Fig. 3. Specifically, in our experiment, we set $C_{0}=768$ and $C_{1}=1024$, respectively. Further, to improve the text-shape consistency, we add a fidelity loss to the objective (16), where the $d(T_{f}(\boldsymbol{x}),\boldsymbol{y})=cos^{-1}(T_{f}(\boldsymbol{x}),\boldsymbol{y})$ is the spherical distance between the output and the paired image feature. $\lambda$ is the weight, which was set to 1 in our experiment. Finally, the desired $f$ and $g$ pair is obtained by a training procedure with alternate optimization: i) Potential $f$ is optimized by Eq. (16); ii) Potential $g$ is optimized by Eq. (17).
@@ 3.3 Text-to-3D generation with SOT map
After we get the well-optimized potential $f$, the SOT map $T_{f}(\boldsymbol{x})$ from text feature hyper-sphere $\boldsymbol{X}\subset\mathbb{S}^{767}$ to image feature hyper-sphere $\boldsymbol{Y}\subset\mathbb{S}^{767}$, can be obtained with Eq. (15) to achieve semantic alignment. For input texts, we utilize CLIP to obtain the initial embedded text features. In previous approaches, text features were directly used as image features for generation. To bridge the gap between the image and text features, we apply the obtained powerful SOT map to project text feature ${\boldsymbol{X}}$ to the aligned $\hat{\boldsymbol{Y}}$ in the CLIP embedding space. Apart from cosine similarity, the SOT map also aligns the two modalities from the view of global distribution. Thereafter, as illustrated in Fig. 2, a diffusion-based conditional generator and a controlled NeRF model are followed to transform the output feature to signed distance functions (SDFs) and texture color. To construct a triangle mesh, a differentiable Marching Cubes [71] is adopted.
@@
4 Experiments
Section 3 analyzes the benefits of our proposed text-to-3D generation framework HOTS3D from the theoretical perspective, and then we will further evaluate the performance of HOTS3D experimentally. The experimental results indicate that HOTS3D has the following pros: 1) it serves as a plug-and-play tool for semantic alignment of CLIP-guided models; 2) the text-3D consistency of shape generated by our model has been drastically enhanced.
4.1 Setup
4.1.1 Datasets
The text prompts are from dataset Text2Shape [15] which provides descriptions for the ‘chair’ and ‘table’ categories in ShapeNet. These text prompts we utilized for testing are the first 1800 text prompts of Text2Shape test dataset. Specifically, we synthesize 2 shapes for each prompt from 1800 text prompts, then, employ Blender [72] to render 20 images for each synthesized 3D mesh, which were used for metrics calculation in subsequent experiments. The ground truth images are obtained by choosing the corresponding mesh from ShapeNet and rendering 20 views for each shape. All image sizes are uniformly set to $512\times 512$. When rendering the view, we set up simple lighting and materials: all models were rendered using a fixed lighting configuration that only supports diffuse and ambient shading. In Fig. 4, several exemplary text-image pairs are shown. For simplicity, each shape only displays an image from one view.
Figure 4: Example text-image pairs of Text2Shape dataset.
4.1.2 Baseline Methods
We compare our method with six state-of-the-art methods, including four end-to-end generative models: CLIP-Forge [2], Shap·E [18], SDFusion [8], TAPS3D [33], as well as two optimization based methods: PureCLIPNeRF [3] and CLIP-Mesh [19]. For a fair comparison, we conduct tests using the default settings, official implementations, and pre-training models for all baseline methods.
Specifically, CLIP-Forge is a zero-shot text-to-shape generative model that addresses the scarcity of paired data. Shap·E, a latent diffusion model over a space of 3D implicit functions, can render both NeRFs and textured meshes. Due to its extensive training dataset and diverse generation capabilities, Shap·E serves as the backbone model for our HOTS3D. SDFusion employs diffusion models for signed distance functions, facilitating various applications such as 3D shape generation, shape completion, and texture generation. TAPS3D focuses on textured shape generation and is trained with the ShapeNet Table and Chair classes. PureCLIPNeRF optimizes model parameters by predicting the density and color corresponding to voxels, with text-based generation relying solely on CLIP [14]. Finally, CLIP-Mesh generates textured meshes through continuous optimization of the vertex coordinates and corresponding texture maps of the initial mesh.
4.1.3 Evaluation Metrics
@@
For the qualitative evaluation, we employ FID [73] and LFD [74] score to assess the quality of the generated shapes. FID measures the distance between the distribution of the rendered images and the ground truth image distribution. LFD measures the dissimilarity between a groundtruth mesh and a generated mesh with rendered images from 60 views. Both FID and LFD indicate higher quality with lower scores. Moreover, to measure the consistency between the generated shapes and input text, we report the CLIP retrieval precision (CLIP R-Precision) [75] on the test dataset with diverse text prompts and objects. This metric quantifies the percentage of generated images that the CLIP encoder associates with the correct text prompt used for generation. For a fair comparison, we followed CLIP R-Precision [75] to calculate the top-10 retrieval accuracy when retrieving the matching text from 100 candidates using the generated image as a query.
Method CLIP R-Precision $\uparrow$ FID $\downarrow$ LFD $\downarrow$ S-A100
@@
SDFusion [8] 72.3 72.8 173.8 5929 8.9
TAPS3D [33] 66.6 67.6 139.2 7530 9.7
Ours 79.3 80.1 42.6 5349 10.2
PureCLIPNeRF [3] 82.8 81.0 172.3 - 1597.9
CLIP-Mesh [19] - - - - 7366.9
TABLE I: Quantitative comparison on Text-to-3D generation. Two distinct models are utilized to compute the CLIP R-Precision. S-A100 denotes the average time (in seconds) it takes to generate 20 images of a shape for a text on the NVIDIA Tesla A100 GPU. Methods of the last two rows are CLIP-guided optimization-based.
CLIP-Forge    PureCLIPNeRF    TAPS3D       SDFusion      Shap·E       Ours
@@
Figure 5: Qualitative comparison. The corresponding text is displayed above each row and the text in the red box is the corresponding area for distinctive description.
4.2 Experiment Results
4.2.1 Implementation Details
We only need to train SOT for the whole network, while other modules remain frozen. Specifically, we utilize the CLIP ViT-L/14 encoder to extract text features as input and image features of rendered shapes as ground truth for SOT training with objectives defined in Eq. (15) and Eq. (16) of our main manuscript. The weight $\lambda$ in Eq. (15) was empirically set to 1 in our experiment. The learning rate is $10^{-4}$, and the optimizer is Adam [76]. Training the SOT map on the dataset Text2Shape [15] takes approximately 28 hours with a Tesla A100 GPU.
@@ 4.2.2 Quantitative Results
We report the CLIP retrieval precision metrics and FID in Tab. I, and both the CLIP ViT-B/16 and CLIP ViT-B/32 encoders are employed as retrieval models. Compared to the first four CLIP-guided end-to-end generative models, our HOTS3D achieves the highest CLIP R-Precision with both retrieval models. Compared to optimization-based methods, our HOTS3D can achieve comparable performance with significantly lower time costs. By leveraging the SOT map, our method bridges the modality gap between text and image features and thus can generate more faithful 3D structures to the input text. Moreover, with the SOT map, we can compensate for the missing detailed information in texts composed of high-level semantics. Consequently, our method exhibits a significantly lower FID score compared to these baseline methods.
To investigate whether semantic alignment can benefit shape quality, we evaluated the dissimilarity between real meshes and generated meshes using the LFD metric, as shown in Tab. I. Since PureCLIPNeRF [3] outputs rendered images from different views instead of meshes, it cannot be evaluated using LFD. Additionally, CLIP-Mesh [19] was excluded from the comparison due to its lengthy generation time. As seen in Tab. [19], our method achieved the best LFD score, indicating that it produces high-quality 3D shapes consistent with the target objects.
We also present a comparison of the generation efficiency of a single text. As shown in the Tab. I, CLIP-guided optimization methods [3, 19] require training the model for each input text, so it often takes tens of minutes to process a single text input well. Especially for CLIP-Mesh, it takes over two hours to process a text. Due to issues of expensive time costs, we abandoned the metric score test for CLIP-Mesh. In comparison, our method can generate a 3D shape that is high-quality and highly consistent with the input text in just about 10 seconds. Compared to Shap·E, the SOT map of our method only incurs an additional expense of 3.2 A100 seconds (See Tab. I) which is marginal for 3D generation tasks, but improves substantially in all metrics, especially in FID score.
@@
Input $\boldsymbol{X}$ Distribution     Target $\boldsymbol{Y}$ Distribution      Spherical Coordinates  Euclidean Coordinates
(a) *
Figure 6: Visualization of two SOT map results. The first two columns indicate distributions of input and target, respectively (warmer color means higher probability). The last two columns are the visualization of the input point set $\boldsymbol{X}$, target point set $\boldsymbol{Y}$, and mapped results of SOT $T(\boldsymbol{X})$ by Mollweide projection and vertical projection, respectively. More overlap between pink SOT mapped results and blue target indicates higher accuracy. Here, the distribution of $T(\boldsymbol{X})$ is highly consistent with $\boldsymbol{Y}$, and $T(\boldsymbol{X})$ effectively covers $\boldsymbol{Y}$ well (overlap of pink and blue).

### v2

@@ 3.2 Derivation of SOT solution on hyper-sphere
Here $\mathtt{CVX}(\nu)$ denotes the set of all convex functions in $L^{1}(\nu)$, $g^{c}(\boldsymbol{x})=\inf_{\boldsymbol{y}}(\log(2-\langle\boldsymbol{x},\boldsymbol{y}\rangle)-g(\boldsymbol{y}))$ is the conjugate c-transform of $g(\cdot)$. According to [84, Theorem 5.26], any optimal pair potential $(g,g^{c})$ in Eq. (6) induce an OT map pushing forward $\nu_{S}$ onto $\mu_{S}$, and $(f,f^{c})$ induce an OT map pushing forward $\mu_{S}$ onto $\nu_{S}$. Specifically, if $c(\boldsymbol{x},\boldsymbol{y})=\frac{1}{2}||\boldsymbol{x}-\boldsymbol{y}||^{2}$, the OT map $T(\boldsymbol{x})=\boldsymbol{x}-\nabla f(\boldsymbol{x})$. When we shift our attention to hyper-spheres, this solution no longer applies. To provide a functional relationship between the OT map and Kontarovich’s potential on hyper-spheres, we introduce some theorems below.
@@
([84, Villani]). Given $\mu$ and $\nu$ on a compact convex domain $\Omega\subset\mathbb{R}^{d}$, there exists an OT plan for the cost $c(\boldsymbol{x},\boldsymbol{y})=h(\boldsymbol{x}-\boldsymbol{y})$, with $h$ strictly convex. It is unique and of the form $(id,T_{\#})\mu$ (id: identity map), provided that $\mu$ is absolutely continuous and $\partial\Omega$ is negligible. Moreover, there exists a Kantorovich’s potential $\varphi$, and the OT map $T$ can be represented as follows:
@@
This theorem was first explained by Villani [84] and later organized by Lei et al. [85, 86]. Based on it, we further provide an explicit functional relationship between the OT map and Kantorovich’s potential $\varphi$ for hyper-spheres case.
@@
Given two probability measures $\mu_{S}$ and $\nu_{S}$ which are defined on $\boldsymbol{X}\subset\mathbb{S}^{d}$ and $\boldsymbol{Y}\subset\mathbb{S}^{d}$, we have
@@
Proposition 3.3 transforms the solution of the SOT map into the problem of optimizing the $W_{d}$ with $f$ and $g$ as variables. Its advantage lies in establishing a direct connection between the $f$ and $g$ and transforming the infimum calculation of c-transform into a parameter optimization process during the learning. Moreover, the gradient of potential $f$ induces the SOT map, which can fit discontinuous maps and allow sharp boundaries to be kept.
According to Theorem 3.2, there exists a potential $f$ such that $T_{f}:\boldsymbol{X}\to\boldsymbol{Y}$ in Eq. (15) satisfies (4). This indicates that there exists an optimal pair $(f_{0},g_{0})$ achieving the infimum and supremum respectively, and $T_{f_{0}}$ is the OT map from $\mu_{S}$ to $\nu_{S}$. In other words, $T_{f_{0}}$ maps $\boldsymbol{X}$ to $\boldsymbol{Y}$ in a measure preserving manner, satisfying $T(\boldsymbol{X})=\boldsymbol{Y}$. Thus, the solution for Eq. (6) is also a solution for Eq. (14). Based on this, the next pertinent question is how to construct a suitable $f$ and $g$ and effectively solve it in the text-to-3D generation task.
Input $\boldsymbol{X}$ Distribution Target $\boldsymbol{Y}$ Distribution    M & V projection of SSW [77]    M & V projection of Ours
(a) *
Figure 5: Visualization of two SOT map results. The first two columns indicate distributions of input and target (warmer color means higher probability). The last four columns are the visualization of the input point set $\boldsymbol{X}$, target point set $\boldsymbol{Y}$, and mapped results of SSW [77] $SSW(\boldsymbol{X})$ and our SOT $T(\boldsymbol{X})$ by Mollweide (M) projection and vertical (V) projection. More overlap between pink SOT mapped results and blue target indicates higher accuracy. Here, the distribution of $T(\boldsymbol{X})$ is highly consistent with $\boldsymbol{Y}$ (overlap of pink and blue) and maintains clear boundaries.
3.3 Text-to-3D generation with SOT map
Inspired by Makkuva et al. [87], we instantiated the Kantorovich potential $f$ and $g$ with ICNN [26], where ICNNs are a specific parametric class of convex functions with rich enough representation. The network architecture of potential $f$ and $g$ is shown in Fig. 4. Specifically, in our experiment, we set $C_{0}=768$ and $C_{1}=1024$, respectively. Further, to improve the text-shape consistency, we add a fidelity loss to the objective (16), where the $d(T_{f}(\boldsymbol{x}),\boldsymbol{y})=cos^{-1}(T_{f}(\boldsymbol{x}),\boldsymbol{y})$ is the spherical distance between the output and the paired image feature. $\lambda$ is the weight, which was set to 1 in our experiment. Finally, the desired $f$ and $g$ pair is obtained by a training procedure with alternate optimization: i) Potential $f$ is optimized by Eq. (16); ii) Potential $g$ is optimized by Eq. (17).
@@ 3.3 Text-to-3D generation with SOT map
After getting the well-optimized $f$, the SOT map $T_{f}(\boldsymbol{x})$ from text feature hyper-sphere $\boldsymbol{X}\subset\mathbb{S}^{767}$ to image feature hyper-sphere $\boldsymbol{Y}\subset\mathbb{S}^{767}$, can be obtained with Eq. (15) to achieve semantic alignment. For input texts, we utilize CLIP to obtain the initial embedded text features. In previous approaches, text features were directly used as image features for generation. To bridge the gap between the image and text features, we apply the obtained SOT map to project text feature ${\boldsymbol{X}}$ to the aligned $\hat{\boldsymbol{Y}}$ in the CLIP embedding space. Apart from spherical distance, the SOT map also aligns the two modalities from the view of global distribution. Thereafter, as illustrated in Fig. 3, a conditional generator is followed to transform the output feature to signed distance functions (SDFs) and texture color. To construct a triangle mesh, a differentiable Marching Cubes [88] is adopted.
4 Experiments
We conduct comprehensive experiments to evaluate the performance of our model. First, we demonstrate the effectiveness of our SOT module, which forms the core of HOTS3D, as described in Sec. 4.1. Next, we showcase HOTS3D’s capability in text-to-3D generation using the Text2Shape [14] and Objaverse [89] datasets, as detailed in Sec. 4.2. Finally, we perform ablation studies to analyze the impact of our SOT module on 3D generation in  Sec. 4.2.5.
4.1 Effectiveness of the SOT Map
To demonstrate the effectiveness of the proposed SOT algorithm in section 3.2, we conducted comparative experiments with SSW [77] on synthesized spherical data, and the visualization results are shown in Fig. 5.
Here, we employed Mollweide projection [90] and vertical projection to visualize the results. The Mollweide projection is an area-preserving [91] projection generally used for maps of the world or celestial sphere. It trades the accuracy of angle and shape for the accuracy of proportions in area, making it useful for maps depicting global distributions. Specifically, in the first three and fifth columns of Fig. 5, we used Mollweide projection, and in the fourth and last columns, we employed vertical projection to provide different viewing perspectives. In the first row of Fig. 5, the source domain data $\boldsymbol{X}$ follows four modes Von Mises-Fisher (vMF[92] distribution, while the target data $\boldsymbol{Y}$ follows uniform distribution on two discontinuous one-eighth spheres. In the second row, the source $\boldsymbol{X}$ follows single mode vMF and the target $\boldsymbol{Y}$ follows six modes vMF.
The vMF distribution, often regarded as the spherical counterpart to the Gaussian distribution, is a distribution on $S^{d-1}$ characterized by a concentration parameter $\kappa>0$ and a location parameter $\mu\in S^{d-1}$ through the density
$\forall\theta\in S^{d-1},\ f_{\mathrm{vMF}}(\theta;\mu,\kappa)=\frac{\kappa^{d/2-1}}{(2\pi)^{d/2}I_{d/2-1}(\kappa)}\exp(\kappa\mu^{T}\theta).$
Where $I_{\nu}(\kappa)=\frac{1}{2\pi}\int_{0}^{\pi}\exp(\kappa\cos(\theta))\cos(\nu\theta)\mathrm{d}\theta$ is the modified Bessel function. Specifically, we set the mean direction ($\mu$ in Euclidean space) to $(1,0,0)$, $(0,1,0)$, $(-1,0,0)$ and $(0,-1,0)$ respectively, and then adopt commonly used open source code [^1] https://github.com/dlwhittenbury/von-Mises-Fisher-Sampling/ to generate a mixed vMF distribution with 4 modes. For the target data $Y$, we uniformly sampled on two discontinuous one-eighth spheres, whose elevation $\theta$ and azimuth $\varphi$ ranges are $\frac{\pi}{3}\leq\theta\leq\frac{5\pi}{6},0\leq\varphi\leq\frac{\pi}{2}$ and $\frac{\pi}{3}\leq\theta\leq\frac{5\pi}{6},\frac{\pi}{2}+\frac{\pi}{12}\leq\varphi\leq\pi+\frac{\pi}{12}$. For another experiment, we set the mean direction to $(1,0,0)$, $(0,1,0)$, $(0,0,1)$, $(-1,0,0)$, $(0,-1,0)$, and $(0,0,-1)$ to generate a mixed vMF distribution with 6 modes, and take it as the target distribution. Then, we set the mean direction as $(1,0,0)$ to generate a single mode vMF and take it as the source distribution.
In this experiment, we instantiated the $f$ and $g$ in Fig. 4 with a 4-layer ICNN [26], including 2 Convex Block. We use Adam [93] with learning rate in $10^{-4}$ to optimize the objective of (16) and (17). Here, $C_{0}=3$, $C_{1}=64$, $\lambda=0$.
The results show that, compared to SSW, our SOT map not only aligns the target distribution more accurately but also preserves the boundaries more effectively. This demonstrates the superiority of our method in handling spherical distribution alignment, particularly in the presence of discontinuities in the data distribution.
4.2 Text-to-3d Generation
4.2.1 Setup
Datasets. To benchmark HOTS3D against prior methods, we use the widely adopted ShapeNet [94] and Objaverse [89] datasets. For ShapeNet, we follow the setup in baselines [7, 17, 16] and choose two specific categories ‘chair’ and ‘table’, and the corresponding text prompts are from dataset Text2Shape [14]. The ground truth images, including those used for training and evaluation, are generated by selecting the corresponding meshes from ShapeNet and rendering 20 views for each shape. We employ Blender [95] to render with a simple lighting and material setup: all models were rendered with a fixed lighting configuration that supports only diffuse and ambient shading. We use the first 1800 text prompts of the Text2Shape test dataset for testing.
For Objaverse, we use the paired image-text dataset from [96] to train and test our model. Specifically, they employ Blender [95] to render 20 images for each 3D mesh and BLIP2 [97] for captioning. All image sizes are uniformly set to $512\times 512$. The text prompts we utilized for testing are 2000 texts selected at equal intervals from the entire dataset.
@@ 4.2.1 Setup
More specifically, the generator is implemented as the composite of a conditional diffusion model and a decoder. The conditional diffusion model is with a MLP+Transformer+ MLP architecture: the first MLP integrates the conditional CLIP embedding, and the second MLP processes the implicit field. The NeRF-based decoder is then used to render the implicit field and extract the SDF for mesh generation. The potential $f$ and $g$ are optimized using the objectives defined in (16) and (17), with the weight $\lambda$ empirically set to 1. We use a learning rate of $10^{-4}$ and the Adam optimizer [93]. Training the SOT map on the Text2Shape dataset [14] takes approximately 28 hours on a Tesla A100 GPU.
Baseline Methods. We compare HOTS3D with thirteen baseline approaches, including end-to-end models: SDFusion [7], Shap·E [18], TAPS3D [33], EXIM [39], XCube [38], Michelangelo [40], Ln3diff [58], Kiss3DGen [55], and Hunyuan3D [59], and optimization-based methods: 3DFuse [51], PCLIPNeRF [2], ProlificDreamer [9], RichDreamer [48]. For fair comparisons, we use the default settings, official implementations, and pretrained models for all baseline methods.
@@
Evaluation Metrics. For a comprehensive quantitative evaluation, we employ FID [98], P-FID [99], LFD [100] score, Chamfer Distance (CD), and F-Score to assess the quality of the generated shapes. FID measures the distance between the rendered image distribution and ground truth distribution. LFD quantifies the dissimilarity between a ground truth mesh and a generated mesh by comparing rendered images from 60 viewpoints, while CD measures the distance between point clouds sampled from the two shapes. P-FID is a modified version of FID tailored for point clouds, designed to measure the distance between the distribution of generated point clouds and that of real point clouds. F-Score evaluates the similarity between the generated and the ground-truth surface. Moreover, to measure the consistency between the generated shapes and input text, we report the CLIP R-Precision [101]. This metric quantifies the percentage of generated images the CLIP encoder associates with the correct text prompt used for generation. We followed the original settings of CLIP R-Precision [101] to calculate the top-10 retrieval accuracy when retrieving the matching text from 100 candidates using the generated image as a query. For 3D metrics P-FID, CD, and F-Score, we uniformly sample 4,096 points from meshes.
TABLE II: Quantitative comparison on Text2Shape. Two distinct models are utilized to compute the CLIP R-Precision. S-A100 denotes the average time (in seconds) it takes to process a text on the NVIDIA Tesla A100 GPU. The methods of the last two rows are optimization-based.
Method CLIP R-Precision $\uparrow$ FID $\downarrow$ LFD $\downarrow$ S-A100$\downarrow$
@@
SDFusion 0.723 0.728 173.8 5929 8.9
TAPS3D 0.666 0.676 79.2 6530 9.7
EXIM 0.714 0.721 45.3 6005 21.7
Michelangelo 0.746 0.752 100.9 6632 7.8
Ours 0.793 0.801 42.6 5349 10.2
PCLIPNeRF 0.828 0.810 172.3 - 1597.9
3DFuse 0.823 0.829 148 - 2417.6
Kiss3DGen   Ln3Diff   TAPS3D   SDFusion  Shap·E  Ours
@@
Figure 6: Qualitative comparison on Text2Shape. Red boxes on results match the distinctive descriptions highlighted in red.
@@ 4.2.2 Quantitative Results
For quantitative comparisons, we report the CLIP retrieval precision and other shape assessment metrics in Tab. II and Tab. III, and both the CLIP ViT-B/16 and CLIP ViT-B/32 encoders are employed as retrieval models. Compared to end-to-end models, our HOTS3D achieves the highest CLIP R Precision with both retrieval models. Compared to optimization based methods, our HOTS3D can achieve comparable performance with significantly lower time costs. By leveraging the SOT module, our method bridges the modality gap and thus can generate more faithful 3D assets to the input text. Moreover, by aligning text features to multiple image features from the same object, we can compensate for the missing detailed information in texts composed of high-level semantics. Consequently, our method exhibits better scores than these baseline methods.
To investigate whether semantic alignment can benefit shape quality, we evaluated the dissimilarity between real meshes and generated meshes using the LFD metric. Since LFD is a mesh-based metric and the final outputs of PCLIPNeRF and 3DFuse are rendered images from multiple views, they cannot be used to compute LFD scores. As seen in  Tab. II, our method achieved the best LFD score, indicating that it produces high-quality 3D shapes consistent with the target objects. We also present a comparison of the generation efficiency of a single text. As shown in the Tab. II, optimization-based methods [2, 51] require training the model for each input text, so it often takes tens of minutes to process a single text well. In comparison, our method can generate 3D shapes that are high-quality and highly consistent with input texts in just about 10 seconds. Compared to Shap·E, the SOT map of our method only incurs an additional expense of 3.2 seconds, which is marginal for 3D generation, but improves substantially in all metrics, especially in FID score. Compared with the SDS-based optimization method 3DFuse [51]. Our HOTS3D demonstrates comparable scores while significantly reducing time costs.
@@
TABLE III: Quantitative comparison on Objaverse.
[…]
ViT-B/32 ViT-L/14
[…]
SDFusion 0.712 0.725 6.8 9447 0.128 0.747
TAPS3D 0.807 0.803 20.2 10644 0.119 0.869
Xcube 0.790 0.796 3.0 7837 0.150 0.957
Michelangelo 0.864 0.883 3.2 10321 0.154 0.893
Ln3diff 0.835 0.862 2.8 7516 0.139 0.928
Hunyuan3D 0.878 0.901 5.7 7671 0.150 0.915
Kiss3DGen 0.800 0.846 7.1 8203 0.142 0.953
Ours 0.900 0.911 2.5 4640 0.068 0.976